OpenStudy (anonymous):

find a linear model and use it to make a prediction. Q: there are 2 leaves along 3in.of and ivy vine. there are 14 leaves along 15in of the same vine. how many leaves are there along 6in.ofthe vine?

OpenStudy (anonymous):

let the length of the vine be x let the number of leaves by y we have two ordered pairs given (3,2) and (15,14) slope = (14-2)/(15-3) = then y = x + b using (3,2) 2 = 3 + b b = -1 then y = x - 1 , where x ≥ 1 so when x = 6, y = 5 there would be 5 leaves for a 6 inch long vine.